# thanks: Corresponding author

### v2

@@ Abstract
We extend previous metric-based Schwarzschild studies of gravitational–electromagnetic (GEM) coupling to rotating black holes by working directly in a curvature-based Newman–Penrose/Teukolsky framework on Kerr spacetime. Within a minimally coupled Einstein–Maxwell system, we derive explicit quadratic electromagnetic source terms for the spin-2-2 Teukolsky equation, providing a foundation for future numerical studies of GEM interactions in the framework of black-hole spectroscopy. Moreover, we give order-of-magnitude arguments showing that GEM quadratic quasinormal modes (QQNMs) can become relevant in a range of charged and magnetized astrophysical scenarios. Finally, we show through a brief dilaton-theory example that the GEM QQNM spectrum is sensitive to how gravity couples to electromagnetism, thereby providing a model-based way to test minimal coupling and to constrain hidden $U(1)$ sectors with gravitational-wave observations.
@@ Introduction
Studying nonlinear effects has become a cornerstone of modern gravitational-wave theory and phenomenology, providing deep insights into the intricate dynamics of strong-field regimes. Numerous studies have presented compelling evidence of nonlinear imprints in the ringdown phase of numerical waveforms simulating binary black hole (BBH) mergers [1, 2, 3, 4, 5, 6]. Perturbative second-order calculations now play a critical role in both ringdown analyses [7, 8, 9, 10, 11, 12, 13, 14, 15] and the modeling of extreme-mass-ratio inspiral (EMRI) waveforms [16, 17, 18, 19, 20]. Nonlinearities are anticipated to have significant implications for the analysis of future high signal-to-noise ratio detections with Advanced LIGO, next-generation ground-based detectors such as the Einstein Telescope and Cosmic Explorer, and the upcoming LISA mission [21, 22, 23, 24, 25, 26, 27, 28, 29], thereby shedding further light on strong-gravity physics.
@@
On the theoretical side, examining GEM effects offers a valuable framework for testing alternative gravity models and the principle of minimal coupling, and for exploring a broader class of applications. Nevertheless, GEM coupling could also contribute to the evolving landscape of multi-messenger astronomy, particularly following the detection of the GW170817 event [34, 31], which provided the first observation of gravitational waves (GWs) and electromagnetic radiation across the electromagnetic spectrum. With advances in observational instruments, many other multi-messenger sources could be observed in the future [35, 36]. Hence, from a phenomenological perspective, a natural step forward is to explore how these fields interact and influence each other. If GWs and electromagnetic radiation can leave mutual imprints on one another, different radiation channels will provide complementary and overlapping information about the same astrophysical event, offering new opportunities to probe and refine our understanding of the strong-field regime.
In this work, we focus on the GEM sector and build on previous analyses of GEM coupling on Schwarzschild backgrounds [37, 38]. We take the next step by extending this framework to the more astrophysically relevant case of rotating black holes. Our primary goal is to explore the structure of GEM interactions on Kerr spacetime, deriving the corresponding quadratic sources for the gravitational Teukolsky equation and casting them in a form suitable for practical applications. In addition, we present several plausibility arguments for the significance of GEM interactions across different astrophysical scenarios, namely neutron star–black hole (NSBH) EMRIs, binary neutron star (BNS) mergers as well as charged BBH mergers. Moreover, we discuss how GEM coupling is theory-dependent and illustrate, through a brief example, how this dependence can be used to test the minimal coupling principle.
The structure of this paper is as follows. In Sec. I we formulate GEM coupling on Kerr spacetime within the Newman–Penrose formalism and derive explicit expressions for the quadratic electromagnetic source of the gravitational Teukolsky equation. In Sec. LABEL:GEM_Coupling we briefly recap the associated GEM QQNMs and their relation to linear gravitational and electromagnetic QNMs. In Sec. III we assess the plausibility of observing GEM effects in several astrophysically motivated scenarios. We summarize our findings and outline future directions in Sec. IV.
@@ I GEM in Kerr spacetime
According to the procedure outlined in [38], analyzing GEM coupling for the gravitational sector involves first solving for the electromagnetic perturbation $\phi^{(1)}_{i}$ (with $i=0,1,2$) and then using the resulting electromagnetic stress-energy tensor as input for the gravitational perturbation equations. The specific calculations may vary depending on the scenario. For EMRIs, the first-order electromagnetic perturbation equations are sourced by a non-zero four-current in addition to the initial conditions. On the other hand, for ringdown analysis, the remnant is treated as a perturbed stationary background (e.g., a Kerr black hole in this work), where only initial data is relevant.
In either case, the particular solution is obtained through convoluting the Green’s function—for the Teukolsky kernel [18]—of the electromagnetic perturbation with the effective source. The resulting solution is then used to construct the effective source for the gravitational perturbation via the electromagnetic stress-energy tensor [38, 37], allowing for the analysis of GEM modes at the first level of mixing [^1] The first level of mixing refers to the hierarchy between the energy scales of “pure” gravitational phenomena and electromagnetic ones. See [38, 37] for further discussion.. This is a concise sketch of the procedure. Here we only focus on the perturbed, minimally coupled Einstein–Maxwell system on the Kerr background.
@@
As shown in [38], the $T_{4}^{\mathrm{EM}}$ can be rearranged to be decomposed into two pieces: $T^{\text{lin}}_{4}\left(\phi^{(1)}_{i},J^{\text{Matter}}\right)$ and $T^{\text{quad}}_{4}\left(\phi^{(1)}_{i}\phi^{(1)}_{j}\right)$. The first term is primarily relevant in EMRI problems and controls GEM contributions that are linear in the electromagnetic perturbations, while the second term is quadratic. As discussed in this work and further analyzed in [38, 37], minimal coupling partially control how GEM effects manifest as linear and quadratic in the electromagnetic perturbations.
@@
As previously mentioned, the source term depends on $\Phi^{(1)}_{i}$, its first derivative $\partial_{\mu}\Phi^{(1)}_{i}$, and various background quantities. Even if the system is not explicitly separable, solving (4) remains computationally manageable. However, constructing $\Phi_{1}$ can be somewhat more challenging, though it is still far less computationally intensive than reconstructing the metric for second-order “pure” gravitational perturbations.
@@ I.2 Gravity-Source Separation
The particular solution to the cases mentioned earlier are expected to take a separable form. A practical approach involves leveraging the orthonormality of spin-weighted spheroidal harmonics [41].
@@
However, terms involving $\frac{1}{\rho^{*p}}$ where $p=1,2,3$, can complicate the integration. Under the integration, these terms necessitate numerical evaluation on a 2D grid of $(r,\theta)$, which might be computationally expensive. Otherwise, the calculation is simpler and can be performed either numerically by integrating only over $\theta$ or analytically by expanding each spin-weighted spheroidal harmonic $S_{\ell m}$ in terms of spin-weighted spherical harmonics $Y_{\ell m}$ and use Clebsch-Gordan coefficients (3j-symbols) under nested sums.
@@
where $f_{l}(r)$ encodes the radial dependence. This expansion introduces an additional summation over $l$, but it simplifies the problem by decoupling the angular and radial components. Higher powers of $\frac{1}{\rho^{*}}$ can be obtained recursively by taking derivatives of (14). The resulting series should resemble known multipole expansions in electrodynamics textbooks.

### v3

@@ Abstract
We extend our previous Schwarzschild metric-based studies of gravitational–electromagnetic (GEM) coupling to rotating black holes by working directly in a curvature-based Newman–Penrose/Teukolsky framework on Kerr spacetime. Within a minimally coupled Einstein–Maxwell system, we derive explicit quadratic electromagnetic source terms for the spin-2-2 Teukolsky equation, providing a foundation for future numerical studies of GEM interactions in the framework of black-hole spectroscopy. Moreover, we give order-of-magnitude arguments showing that GEM quadratic quasinormal modes (QQNMs) can become relevant in a range of charged and magnetized astrophysical scenarios. Finally, we show through a brief dilaton-theory example that the GEM QQNM spectrum is sensitive to how gravity couples to electromagnetism, thereby providing a model-based way to test minimal coupling and to constrain hidden $U(1)$ sectors with gravitational-wave observations.
@@ Introduction
Studies of nonlinear effects have become a focus of the recent progress in gravitational-wave theory and phenomenology, providing deep insights into the intricate dynamics of strong-field regimes. Numerous studies have presented compelling evidence of nonlinear imprints in the ringdown phase of numerical waveforms simulating binary black hole (BBH) mergers [1, 2, 3, 4, 5, 6]. Perturbative second-order calculations now play a critical role in both ringdown analyses [7, 8, 9, 10, 11, 12, 13, 14, 15] and modeling of extreme-mass-ratio inspiral (EMRI) waveforms [16, 17, 18, 19, 20]. Nonlinearities are anticipated to have significant implications for the analysis of future high signal-to-noise ratio detections with Advanced LIGO, next-generation ground-based detectors such as the Einstein Telescope and Cosmic Explorer, and the upcoming LISA mission [21, 22, 23, 24, 25, 26, 27, 28, 29], thereby shedding further light on strong-gravity physics.
@@
On the theoretical side, examining GEM effects offers a valuable framework for testing alternative gravity models and the principle of minimal coupling, and for exploring a broader class of applications. Nevertheless, GEM coupling could also contribute to the evolving landscape of multi-messenger astronomy, particularly following the detection of the GW170817 event [34, 31], which provided the first observation of gravitational waves (GWs) and electromagnetic radiation across the electromagnetic spectrum. With advances in observational instruments, many other multi-messenger sources could be observed in the future [35, 36]. Hence, from a phenomenological perspective, a natural step forward is to explore how these fields interact and influence each other. If GWs and electromagnetic radiation can leave mutual imprints on each other, different radiation channels will provide complementary and overlapping information about the same astrophysical event, offering new opportunities to probe and refine our understanding of the strong-field regime.
In this work, we focus on the GEM sector and build on previous analysis of GEM coupling on Schwarzschild backgrounds [37, 38]. We take the next step by extending this framework to the more astrophysically relevant case of rotating black holes. Our primary goal is to explore the structure of GEM interactions on Kerr spacetime, deriving the corresponding quadratic sources for the gravitational Teukolsky equation and casting them in a form suitable for practical applications. In addition, we present several plausibility arguments for the significance of GEM interactions across different astrophysical scenarios, namely neutron star–black hole (NSBH) EMRIs, binary neutron star (BNS) mergers as well as charged BBH mergers. Moreover, we discuss how GEM coupling is theory-dependent and illustrate, through a brief example, how this dependence can be used to test the minimal coupling principle.
The structure of this paper is as follows. In Sec. I we formulate GEM coupling on Kerr spacetime within the Newman–Penrose formalism and derive explicit expressions for the quadratic electromagnetic source of the gravitational Teukolsky equation. In Sec. II we briefly recap the associated GEM QQNMs and their relation to linear gravitational and electromagnetic QNMs. In Sec. III we assess the plausibility of observing GEM effects in several astrophysically motivated scenarios. We summarize our findings and outline future directions in Sec. IV.
@@ I GEM in Kerr spacetime
According to the procedure outlined in [38], analyzing the GEM coupling for the gravitational sector involves first solving for the electromagnetic perturbation $\phi^{(1)}_{i}$ (with $i=0,1,2$) and then using the resulting electromagnetic stress-energy tensor as input for the gravitational perturbations equations. Specific calculations may vary depending on the scenario. For EMRIs, the first-order electromagnetic perturbation equations are sourced by a non-zero four-current in addition to the initial conditions. On the other hand, for ringdown analysis, the remnant is treated as a perturbed stationary background (e.g., a Kerr black hole in this work), where only initial data is relevant.
In either case, the particular solution is obtained through convoluting the Green’s function (for the Teukolsky kernel [18]) of the electromagnetic perturbation with the effective source. The resulting solution is then used to construct the effective source for the gravitational perturbation via the electromagnetic stress-energy tensor [38, 37], allowing for the analysis of GEM modes at the first level of mixing [^1] The first level of mixing refers to the hierarchy between the energy scales of “pure” gravitational phenomena and electromagnetic ones. See [38, 37] for further discussion.. This is a concise sketch of the procedure. Here we only focus on the perturbed, minimally coupled Einstein–Maxwell system on the Kerr background.
@@
As shown in [38], $T_{4}^{\mathrm{EM}}$ can be rearranged to decompose into two pieces: $T^{\text{lin}}_{4}\left(\phi^{(1)}_{i},J^{\text{Matter}}\right)$ and $T^{\text{quad}}_{4}\left(\phi^{(1)}_{i}\phi^{(1)}_{j}\right)$. The first term is primarily relevant in EMRI problems and controls GEM contributions that are linear in the electromagnetic perturbations, while the second term is quadratic. As discussed in this work and further analyzed in [38, 37], minimal coupling partially controls how GEM effects manifest as linear and quadratic in electromagnetic perturbations.
@@
As previously mentioned, the source term depends on $\Phi^{(1)}_{i}$, its first derivative $\partial_{\mu}\Phi^{(1)}_{i}$, and various background quantities. Even if the system is not explicitly separable, solving (4) remains computationally manageable. However, constructing $\Phi_{1}$ can be somewhat more challenging, though it is still far less computationally intensive than reconstructing the metric for the second-order “pure” gravitational perturbations.
@@ I.2 Gravity-Source Separation
The particular solutions to the cases mentioned earlier are expected to take a separable form. A practical approach involves leveraging the orthonormality of spin-weighted spheroidal harmonics [41].
@@
However, terms involving $\frac{1}{\rho^{*p}}$ where $p=1,2,3$, can complicate integration. Under the integration, these terms require a numerical evaluation on a 2D grid of $(r,\theta)$, which might be computationally expensive. Otherwise, the calculation is simpler and can be performed either numerically by integrating only over $\theta$ or analytically by expanding each spin-weighted spheroidal harmonic $S_{\ell m}$ in terms of spin-weighted spherical harmonics $Y_{\ell m}$ and use Clebsch-Gordan coefficients (3j-symbols) under nested sums.
@@
where $f_{l}(r)$ encodes the radial dependence. This expansion introduces an additional summation over $l$, but it simplifies the problem by decoupling the angular and radial components. Higher powers of $\frac{1}{\rho^{*}}$ can be obtained recursively by taking derivatives of (14). The resulting series should resemble the known multipole expansions in electrodynamics textbooks.